# MetaGFN: Exploring Distant Modes with Adapted Metadynamics for Continuous GFlowNets

### v1

@@ 7   $z_{t}\leftarrow z(x_{t})$
9     $\hat{N}\leftarrow\hat{N}+\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
10     $\hat{R}\leftarrow\hat{R}+r(x_{t})\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
@@ 11     $\hat{V}\leftarrow-\frac{1}{\beta}\log\left(\hat{R}/(\hat{N}+\epsilon)+\epsilon\right)$
12     $V_{\text{bias}}(z)\leftarrow V_{\text{bias}}(z)+n\cdot\Delta t\cdot w\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
@@
  propagate $x_{t},p_{t}$ by $\Delta t$ using Langevin dynamics with computed force $F$ (Alg. 2, Appendix D).
@@
Training GFlowNets with Adaptive Metadynamics (MetaGFN) Each Adaptive Metadynamics sample $x_{i}\in\mathcal{X}$ is an off-policy terminal state sample. To train a GFlowNet, complete trajectories are required. We generate these by backward sampling from the terminal state, giving a trajectory $\tau=(s_{0},s_{1},\ldots,s_{n}=x_{i})$, where each state $s_{i-1}$ is sampled from the current backward policy distribution $\hat{p}_{B}(s_{i-1}|s_{i};\theta)$, for $i$ from $n$ to $1$. This approach means that the generated trajectory $\tau$ has reasonable credit according to the loss function, thereby providing a useful learning signal. However, since this requires a backward policy, this is compatible with DB, TB, and STB losses, but not FM loss. Given the superior credit assignment of the former losses, this is not a limitation Madan et al. (2022).
Additionally, we use a replay buffer. Due to the theoretical guarantee that Adaptive Metadynamics will eventually sample all collective variable space (Theorem 3.1), AM samples are ideal candidates for storing in a replay buffer. When storing these trajectories in the replay buffer, there are two obvious choices:
@@
    Store only the Adaptive Metadynamics sample and regenerate trajectories using the current backward policy when retrieving from the replay buffer.
We investigate both options in our experiments. We call the overall training algorithm MetaGFN, with pseudocode presented in Algorithm 3, Appendix D.
@@ 4 Experiments
4.1 Line environment
Figure 1: Line Environment reward function, equation (2).
We consider a one-dimensional line environment, with state space $\mathcal{S}=\mathbb{R}\times\left\{t\in\mathbb{N},1\leq t\leq 3\right\}$, where $t$ indexes the position of a state in a trajectory. The source state is $s_{0}=(0,0)$ and trajectories terminate after exactly $3$ steps. The terminal states are therefore $\mathcal{X}=\mathbb{R}\times\left\{3\right\}\cong\mathbb{R}$. The reward density, plotted in Figure 1, consists of an asymmetric bimodal peak near the origin and an additional distant lone peak. It is given by the Gaussian mixture distribution:
@@ 4.1 Line environment
where $\mathcal{N}(\mu,\sigma^{2})$ is a Gaussian density with mean $\mu$ and variance $\sigma^{2}$. The forward and backward probability transition kernels are a mixture of $3$ Gaussian distributions which, along with the flow $\hat{f}$, are parameterised by an MLP.
We compare the following exploration strategies: entirely on-policy (no exploration), noisy exploration, Thompson Sampling, and MetaGFN. For each strategy, we use a replay buffer and alternate between an exploration batch and a replay buffer batch. For MetaGFN, we use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$ and always regenerate MetaGFN training trajectories using the current backward policy. As the environment is one-dimensional, the collective variable is simply $z(x)=x$. We evaluate performance by computing the L1 error between the known reward distribution and the empirical on-policy distribution during training (see Appendix E.1 for the full experimental details).
Figure 2: The L1 difference between on-policy and reward distribution during training in the line environment for different loss functions and exploration strategies. The mean is plotted with standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 2. Among the three loss functions, TB loss has the lowest variance loss profiles, and MetaGFN consistently converges to a lower minimum error than all other exploration strategies. Indeed, MetaGFN was the only method that consistently sampled the distant reward peak at $x=20$, while other methods plateau in error after locking onto the central modes (Appendix E.2). Despite this, we observed that in the initial stages of training, all exploration strategies found occasional samples from the distant peak. The reason MetaGFN is the only method that converges is that Adaptive Metadynamics manages to consistently sample the distant peak, even if the on-policy starts to focus on the central modes. This keeps the replay buffer populated with samples from every reward peak during training, which eventually encourages the on-policy to sample from every mode. The small increase in the loss of MetaGFN around batch number $5\times 10^{3}$ happens because the on-policy distribution widens when Adaptive Metadynamics first discovers the distant peak. In Appendix E.2, we show further details of Adaptive Metadynamics in this environment and we compare different MetaGFN variants, with and without noise, and with and without trajectory regeneration. We confirm that the version of MetaGFN presented in Figure 2 (no added noise and always regenerate trajectories) is the most robust variant.
4.2 Grid environment
Figure 3: Grid Environment reward function.
We consider a two-dimensional environment, with state space $\mathcal{S}=\mathbb{R}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$. The source state is $s_{0}=(0,0,0)$ and trajectories terminate after exactly $3$ steps. Terminal states are $\mathcal{X}=\mathbb{R}^{2}\times\{3\}\cong\mathbb{R}^{2}$. The reward function is given by the sum of four symmetric Gaussians with means $\mu=(-7,-7),(-7,7),(7,7),(7,-7)$ and variances $\sigma^{2}=2$ (Figure 3). The forward and backward probability transition kernels are a mixture of $4$ bivariate Gaussian distributions with means and diagonal variance matrices, along with the flow $\hat{f}$, parameterised by an MLP. We consider the same exploration strategies and evaluation measures as for the Line Environment (see Appendix F for full experiment details).
Figure 4: The L1 difference between on-policy and reward distribution during training in the grid environment for different loss functions and exploration strategies. The mean is plotted with standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 4. We see that for most loss functions and exploration strategies (except Thompson sampling with TB loss), the training curves converge to a low loss, indicating that all four modes of the grid environment are well-sampled. However, in all cases, MetaGFN converges faster than other exploration strategies. The fastest convergence occurs for MetaGFN with TB loss; after $10^{3}$ batches MetaGFN has reached a similar loss that noisy exploration reaches after $10^{4}$ batches and on-policy reaches after $\sim 3\times 10^{4}$ batches. Meanwhile, the loss curves have lower error indicating more stable training with MetaGFN in this environment.
4.3 Alanine dipeptide environment
Figure 5: (a) Free energy surface of alanine dipeptide in an explicit solvent with all metastable states annotated. (b) The potential KDE $\hat{V}$ learnt using Adapted Metadynamics with von Mises kernel, concentration $\kappa=10$.
@@ 4.3 Alanine dipeptide environment
One application of continuous GFlowNets is molecular conformation sampling Volokhova et al. (2023). Here, we train a GFlowNet to sample conformational states of alanine dipeptide (AD), a small biomolecule of $23$ atoms that plays a key role in modelling backbone dynamics of proteins Hermans (2011). The metastable states of AD can be distinguished in a two-dimensional CV space of $\phi$ and $\psi$, the two backbone dihedral angles. The resulting free energy surface for AD in explicit water, $V(\phi,\psi)$, obtained after extensive sampling long molecular dynamics simulation is shown in Figure 5(a). The metastable states (energy minima), in increasing energy, are $P_{||}$, $\alpha_{R}$, $C_{5}$, $\alpha^{\prime}$, $\alpha_{L}$ and $\alpha_{D}$.
The state space is $\mathcal{S}=\mathbb{T}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$. The source state is $s_{0}=P_{||}=(-1.2,2.68)$ and trajectories terminate after exactly $3$ steps. Terminal states are $\mathcal{X}=\mathbb{T}^{2}\times\{3\}\cong\mathbb{T}^{2}$. The reward density is given by the Boltzmann weight, $r(\phi,\psi)=\frac{1}{Z}\exp(-\beta V(\phi,\psi))$, where $Z$ is the normalisation constant. The forward and backward probability transition kernels are defined as a mixture of $3$ bivariate von Mises distributions, parameterised through an MLP. We consider the same exploration strategies and evaluation measures as for the Line Environment (see Appendix G for full experimental details).
@@
Figure 6: The L1 difference between on-policy and reward distribution during training in the alanine dipeptide environment for different loss functions and exploration strategies. The mean is plotted with a standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
@@
The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 6. For each loss function, models trained with MetaGFN generally converge to a lower minimum error than all other exploration strategies. For TB loss, however, the average L1 error is marginally higher than on-policy training, but this conceals the fact that the best-case error is smaller. Thus, to better understand this result, we examine the best and worst training runs (as measured by L1 error) for TB on-policy and TB MetaGFN, shown in Figure 7. We see that the best run trained with metadynamics can sample from the rare $\alpha_{L}$ minima, unlike the on-policy run. In the worst run, MetaGFN fails to converge (although this is rare; only one of 10 runs failed). In Table 1, we quantify how often the different AD modes were correctly sampled over the different repeats (a mode is correctly sampled if the on-policy distribution also has a mode within the correct basin of attraction). The only mode not correctly sampled by any method is $\alpha_{D}$, which has a natural abundance approximately $10$ times less frequent than $\alpha_{L}$. We see that TB loss with MetaGFN is the only combination that can consistently sample the majority of modes, whilst noisy exploration and Thompson Sampling both perform worse than on-policy in this environment.

### v2

@@ 7   $z_{t}\leftarrow z(x_{t})$
9     $\hat{N}\leftarrow\hat{N}+\exp\left({-\frac{1}{2}\sum_{i=1}^{d}\frac{(z_{i}-{z}_{i,t})^{2}}{{\sigma_{i}^{2}}}}\right)$
10     $\hat{R}\leftarrow\hat{R}+r(x_{t})\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{d}\frac{(z_{i}-{z}_{i,t})^{2}}{{\sigma_{i}^{2}}}}\right)$
@@ 11     $\hat{V}\leftarrow-\frac{1}{\beta}\log\left(\hat{R}/(\hat{N}+\epsilon)+\epsilon\right)$
12     $V_{\text{bias}}(z)\leftarrow V_{\text{bias}}(z)+n\cdot\Delta t\cdot w\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{d}\frac{(z_{i}-{z}_{i,t})^{2}}{{\sigma_{i}^{2}}}}\right)$
@@
  propagate $x_{t},p_{t}$ by $\Delta t$ using Langevin dynamics with computed force $F$ (Alg. 3, Appendix B).
@@
Metadynamics requires the gradient of the total potential, where $-\nabla V_{\text{total}}(x,t)=-\nabla(V(x)+V_{\text{bias}}(x,t))$. Using the above assumptions, we have $\nabla V(x)=-\nabla r(x)/(\beta^{\prime}r(x))$. However, $r(x)$ is often a computationally expensive black-box function, and its gradient, $\nabla r(x)$, is unknown. While finite differences can estimate $\nabla r(x)$ for smooth, low-dimensional reward distributions, this approach is impractical in high-dimensional spaces. Below, we explain how to adapt metadynamics to this black-box setting and avoid finite difference gradient estimates by storing a dynamically-updated kernel density estimate (KDE) of the potential. We call the modified metadynamics algorithm Adapted Metadynamics (Algorithm 1).
Let $x_{t}$ denote the metadynamics sample at time $t$ and $z(x)=(z_{1}(x),\dots,z_{d}(x))$ be a given set of collective variables, where $z_{i}$ is a one-dimensional coordinate, and $z\in\mathcal{Z}$ is $d$-dimensional. Let $z_{i,t}\vcentcolon=z_{i}(x_{t})$ denote the corresponding $i^{th}$ CV coordinate at time $t$ and assume that the Jacobian $\nabla_{x}z$ is well-defined. We store a discretisation of the KDE marginal potential and bias potentials in CV space $\mathcal{Z}$. Since these are stored in memory, gradient computations are cheap and require no further evaluations of $r(x)$.
Let $\hat{V}(z,t)$ represent the KDE of the marginal potential at time $t$. To compute $\hat{V}(z,t)$, we maintain two separate KDEs: $\hat{N}(z,t)$ for visited states (line 1) and $\hat{R}(z,t)$ for cumulative rewards (line 1). We update these KDEs on-the-fly at the same time the bias potential is updated, which occurs every integer $n$ steps of Langevin dynamics (line 1). If $\mathcal{Z}\cong\mathbb{R}^{d}$, we use Gaussian kernels with kernel width $\sigma=(\sigma_{1},\ldots,\sigma_{k})\in\mathbb{R}^{k}$, matching the width of the Gaussian bias (line 1) [^3] If $\mathcal{Z}\cong\mathbb{T}^{d}$ ($d$-torus), we use von Mises distributions.. This is a reasonable as both are set by the variability length scale of $V(x)$. Finally, the KDE potential $\hat{V}(z,t)$ is then computed as $\hat{V}(z,t)=-\frac{1}{\beta}\log\left(\frac{\hat{R}(z,t)}{\hat{N}(z,t)+\epsilon}+\epsilon\right)$, for a fixed constant $\epsilon>0$ (line 1). We found empirically that $\epsilon$ ensured numerical stability by preventing division by zero and bounding $\hat{V}$ from above. Defining $\hat{V}$ through the ratio $\hat{R}/\hat{N}$ ensures that it rapidly and smoothly adjusts whenever new modes are discovered. Furthermore, we prove that $\hat{V}$ eventually discovers all reward modes in the CV space. More precisely,
Theorem 3.1.
where $V=V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$.
The proof is in Appendix C.
Note that the algorithm can be extended to a batch of trajectories, where each metadynamics trajectory evolves independently, but with a shared $\hat{V}$ and $V_{\text{bias}}$ which receive updates from every trajectory in the batch. We found empirically that this accelerates exploration and reduces stochastic gradient noise during training and this is the version we use in our experiments (Section 4).
3.2 MetaGFN
Each Adapted Metadynamics sample $x_{i}\in\mathcal{X}$ is an off-policy terminal state sample. To train a GFlowNet, complete trajectories are required. We generate these by backward sampling from the terminal state, giving a trajectory $\tau=(s_{0},s_{1},\ldots,s_{n}=x_{i})$, where each state $s_{i-1}$ is sampled from the current backward policy distribution $\hat{p}_{B}(s_{i-1}|s_{i};\theta)$, for $i$ from $n$ to $1$. This approach means that the generated trajectory $\tau$ has reasonable credit according to the loss function, thereby providing a useful learning signal. However, since this requires a backward policy, this is compatible with DB, TB, and STB losses, but not FM loss. Given the superior credit assignment of the former losses, this is not a limitation [15].
Additionally, we use a replay buffer. Due to the theoretical guarantee that Adapted Metadynamics will eventually sample all collective variable space (Theorem 3.1), AM samples are ideal candidates for storing in a replay buffer. When storing these trajectories in the replay buffer, there are two obvious choices:
@@ 3.2 MetaGFN
    Store only the Adapted Metadynamics sample and regenerate trajectories using the current backward policy when retrieving from the replay buffer.
We investigated both options in preliminary experiments (Appendix D.2). Option 2 was found to be uniformly superior and it is the version we use in our main experiments (Section 4). We call the overall training algorithm MetaGFN, with pseudocode presented in Algorithm 2, below.
Input : Forward policy $P_{F}$. Backwards policy $P_{B}$. Loss function $L$.
Parameters : How often to run Adapted Metadynamics batches, freqMD. How often to run replay buffer batches, freqRB. Batch size, $b$. Stride, $n\in\mathbb{Z}^{+}$. Time step, $\Delta t>0$
1
2for each episode do:
3if episode number is divisible by freqMD:
4   Run Adapted Metadynamics (batch size $b$) for time $n\Delta t$, obtain samples $\{x_{1},\dots,x_{b}\}$
5   Push $\{x_{1},\dots,x_{b}\}$ to the replay buffer
7elif episode number is divisible by freqRB:
8   Random sample $\{x_{1},\dots,x_{b}\}$ from the replay buffer
10
11else:
12   Generate trajectories $\{\tau_{1},\dots,\tau_{b}\}$ on-policy
13
14 Compute loss $l=\sum_{i=1}^{b}L(\tau_{i},P_{F},P_{B})$
 Take gradient step on loss $l$
Algorithm 2 MetaGFN
@@ 4 Experiments
We compare MetaGFN with Thompson sampling, noisy, Local Search GFlowNets and nested sampling (Section 2.3). We run experiments in five continuous environments, summarised below. For each exploration strategy, we use a replay buffer and alternate between exploration and replay buffer batches. For MetaGFN, we use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$. Forward and backward kernels are Gaussian/von Mises mixture distributions, with distribution parameters specified at each state by an MLP. In all environments, the additional computational expense of running Adapted Metadynamics was negligible (<5%) compared to the training time of the models. Full experimental details are in Appendix D.
Figure 2: Line environment reward density.
Line environment: A one-dimensional environment with state space $\mathcal{S}=\mathbb{R}\times\left\{t\in\mathbb{N},1\leq t\leq 3\right\}$, where $t$ indexes the position of a state in a trajectory. The source state is $s_{0}=(0,0)$. The terminal states are therefore $\mathcal{X}=\mathbb{R}\times\left\{3\right\}\cong\mathbb{R}$. The collective variable is the identity, i.e. $z=x$. The reward density, plotted in Figure 2, consists of an asymmetric bimodal peak near the origin and an additional distant lone peak. It is given by the Gaussian mixture distribution:
@@
where $\mathcal{N}(\mu,\sigma^{2})$ is a Gaussian density with mean $\mu$ and variance $\sigma^{2}$.
Figure 3: Free energy surface of alanine dipeptide.
@@
Alanine dipeptide environment: One application of continuous GFlowNets is molecular conformation sampling [12]. Here, we train a GFlowNet to sample conformational states of alanine dipeptide (AD), a a small biomolecule of 23 atoms that plays a key role in modelling the dynamics of proteins [44]. The metastable states of AD are distinguished in a 2D CV space defined by the backbone dihedral angles $\phi$ and $\psi$. The resulting free energy surface $V(\phi,\psi)$ in explicit water, obtained after extensive sampling with long molecular dynamics simulations, is shown in Figure 3. The metastable states, in increasing energy, are $P_{||}$, $\alpha_{R}$, $C_{5}$, $\alpha^{\prime}$, $\alpha_{L}$, and $\alpha_{D}$. The state space is $\mathcal{S}=\mathbb{T}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$, with source state $s_{0}=P_{||}=(-1.2,2.68)$. Terminal states are $\mathcal{X}=\mathbb{T}^{2}\times\{3\}\cong\mathbb{T}^{2}$. The reward function is the Boltzmann weight, $r(\phi,\psi)=\frac{1}{Z}\exp(-\beta V(\phi,\psi))$, where $Z$ is the normalisation constant.
Figure 4: Grid environment reward density in dimension 2.
Grid environments: We consider (hyper)grids in $d=2$, $3$ and $4$ dimensions. The state space is $\mathcal{S}=[-15,15]^{d}\times\{t\in\mathbb{N},1\leq t\leq 3\}$. The $k$-dimensional (hyper)grid consists of $2^{k}$ modes, located at the corners of the (hyper)cube $[-10,10]^{d}$. The (hyper)grid is centered on the origin with an edge width of $20$. The reward modes are Gaussians with variance $\sigma^{2}=2$ (Figure 4). We use trajectory lengths of 3, 5, and 6 for dimensions 2, 3, and 4 respectively. The collective variable is the identity, i.e. $z=x$. The source state is the origin and terminal states are in $\mathcal{X}\cong\mathbb{R}^{k}$.
4.1 Results
@@
In each environment, we run experiments with three different loss functions: Detailed Balance (DB), Trajectory Balance (TB) and Subtrajectory Balance (STB). We evaluate performance by computing the L1 error between the known reward distribution and the empirical on-policy distribution during training resulting from $10^{4}$ independent samples. The results, averaged over $10$ repeats, are shown in Figure 5. In 12 out of 15 of the experiments, MetaGFN outperformed all other exploration strategies. We give further analysis below.
Line Environment: Among the losses, TB shows the lowest variance, and for all losses, MetaGFN always converges to a lowest error  [^4] We do not show results for nested sampling on the line environment as the implementation used did not support one-dimensional environments, see Appendix D.. Indeed, MetaGFN is the only method that consistently samples the distant reward peak at $x=20$ (Appendix D.2). Although other strategies occasionally sampled the distant peak, MetaGFN alone converges because it continuously samples this peak, even if the forward policy starts to lock onto the central modes, thus ensuring that the replay buffer is always populated with diverse samples. The slight increase in the loss of MetaGFN around batch number $5\times 10^{3}$ occurs as the on-policy distribution widens when Adapted Metadynamics first discovers the distant peak. Appendix D.2 provides further analysis of Adapted Metadynamics and compares different MetaGFN variants, with and without noise, and with and without trajectory regeneration. We confirm that the variant of MetaGFN presented in Figure 5 (no added noise, always regenerate trajectories) is the most robust.
@@
Alanine Dipeptide Environment: For each loss, MetaGFN generally converges to a lower error than all other exploration strategies. However, for TB loss, the average L1 error is marginally higher than on-policy training, but this conceals the fact that the best-case error is smaller. To better understand this result, we examined the best and worst training runs (as measured by L1 error) for TB on-policy and TB MetaGFN, shown in Figure 6. Note that, unlike on-policy, the best MetaGFN run learns to sample the rare $\alpha_{L}$ mode. In the worst case, MetaGFN fails to converge (although this is rare; only one of 10 runs failed). In Table 1, we quantify how often the different AD modes were sampled over the different repeats (a mode is considered sampled if the on-policy distribution has a mode within the correct basin of attraction). TB loss with MetaGFN is the only combination that consistently samples the majority of modes. The only mode not sampled by any method is $\alpha_{D}$, which has a natural abundance approximately $10$ times less frequent than $\alpha_{L}$.
Figure 5: The L1 difference between on-policy and reward distribution during training for different loss functions and exploration strategies. The mean is plotted with standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
@@
Grid Environments: MetaGFN achieves the lowest error in 7 of 9 experiments, although note that the performance of all methods generally decreases with increasing dimension. This is unsurprising, since the learning task becomes more difficult. For MetaGFN, this is also an expected outcome of the curse of dimensionality of the replay buffer grid; in high-dimensional spaces metadynamics samples encounter high-reward modes less frequently, leading to a replay buffer with less sample diversity. This is a known limitation of metadynamics [26]. Nested Sampling outperforms MetaGFN in one case (Grid 3D, DB loss) but relies on costly MCMC-based pre-training that scales exponentially with dimension and requires reward gradients, unlike MetaGFN. However, the generally favourable performance of MetaGFN could be due to the method discovering an overall larger diversity of off-policy samples, that are refreshed and updated during training, in contrast to the static samples when using Nested Sampling approach.